Circular No. 7726

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN CENTAURUS W. Liller, Viña del Mar, Chile, reports his discovery of a possible nova

(mag 8.6) on two Technical Pan films taken on Oct. 1.010 and 1.011 UT with an orange filter. The new object is located at α = 13^{h}55^{m}.5, δ = −64^{o}16′ (equinox 2000.0). Nothing brighter than mag 11.0 appears at this location on photos taken by Liller on Sept. 19.00. A. C. Gilmore, Mount John Observatory, University of Canterbury, reports the following photoelectric photometry of the new object obtained with the 0.6-m f/16 reflector under marginal conditions (thin cloud, bright moonlight); reference stars were Cousins F218 and F212: Oct. 2.415, V = 9.11, U − B = +0.83, B − V = +1.41, V − R = +0.81, V − I = +1.69, airmass 2.18. Gilmore provides the following position with uncertainty < 10′′ in each coordinate (proper astrometry not possible due to lack of reference stars in the moonlight): α = 13^{h}55^{m}41^{s}.9, δ = −64^{o}15′48′′ (equinox 2000.0). Visual magnitude estimates: Oct. 2.33, 9.2 (A. Jones, Nelson, New Zealand); 2.542, 9.3 (A. Pearce, Nedlands, W. Australia); 2.708, 9.5 (L. Monard, Pretoria, S. Africa); 2.993, 9.4 (A. Amorim, Florianopolis, Brazil); 3.510, 9.6 (Pearce).

SUPERNOVA 2001eo IN UGC 3963 T. Puckett and M. Marcus, Mountain Town, GA, report the discovery

of an apparent supernova (mag 17.5) on an unfiltered CCD frame (limiting mag 19.0) taken with the Puckett Observatory 0.30-m automated patrol telescope on Oct. 1.38 UT. SN 2001eo is located at α = 7^{h}41^{m}31^{s}.10, δ = +51^{o}41′05′′.4 (equinox 2000.0), which is 10′′.0 west and 13′′.5 south of the center of UGC 3963. The new object was confirmed on CCD frames taken on Oct. 2.34 by Marcus (0.60-m reflector). The candidate is not present on images taken by Puckett on 2001 Mar. 10, 18, and 23 (limiting mag ∼ 20.5), or on Palomar Sky Survey images taken on 1999 Mar. 6, 1992 Nov. 25 (limiting mag ∼ 21.0), and 1954 Feb. 25 (limiting mag ∼ 19.7).

COMET C/1996 B4 (SOHO) D. Hammer reports (cf. IAUC 7718) that X. Leprette found a sungraz-

ing comet on SOHO website C3 coronagraph images; reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2001-S70.

1996 UT  α_{2000}  δ_{2000}
Jan. 21.337  20^{h}27.^{m}8  −21^{o}32′
2001 October 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green